Find five rational numbers between. 
-3/2 and 5/3  
Find five rational numbers between. 
-3/2 and 5/3  

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

-3/2 and 5/3
Let us make the denominators same, say 6
i.e., -3/2 and 5/3 can be written as:
-3/2 = (-3 × 3)/(2× 3) = -9/6
5/3=(5×2)/(3×2)=10/6
Five rational numbers between -3/2 and 5/3 = five rational numbers between -9/6 and 10/6
Therefore, Five rational numbers between -9/6 and 10/6 = -1/6, 2/6, 3/6, 4/6, 5/6

Thanks